The sum of two-sevenths of a number and 3 is 9

turn it into a equation.

Answer by stanbon(75887)

The sum of two-sevenths of a number and 3 is 9.
-------
(2/7)x + 3 = 9
(2/7)x = 6
x = 7*6/2
x = 7*3
x = 21
-----------
Cheers,
Stan H.
----------